### v4

# UniF2ace: A Unified Fine-grained Face Understanding and Generation Model

Abstract
Unified multimodal models (UMMs) have emerged as a powerful paradigm in fundamental cross-modality research, demonstrating significant potential in both image understanding and generation. However, existing research in the face domain primarily faces two challenges: (1) fragmentation development, with existing methods failing to unify understanding and generation into a single one, hindering the way to artificial general intelligence. (2) lack of fine-grained facial attributes, which are crucial for high-fidelity applications. To handle those issues, we propose UniF2ace, the first UMM specifically tailored for fine-grained face understanding and generation. First, we introduce a novel theoretical framework with a Dual Discrete Diffusion (D3Diff) loss, unifying masked generative models with discrete score matching diffusion and leading to a more precise approximation of the negative log-likelihood. Moreover, this D3Diff significantly enhances the model’s ability to synthesize high-fidelity facial details aligned with text input. Second, we propose a multi-level grouped Mixture-of-Experts architecture, adaptively incorporating the semantic and identity facial embeddings to complement the attribute forgotten phenomenon in representation evolvement. Finally, to this end, we construct UniF2aceD-1M, a large-scale dataset comprising 130K fine-grained image-caption pairs and 1M visual question-answering pairs, spanning a much wider range of facial attributes than existing datasets. Extensive experiments demonstrate that UniF2ace outperforms existing models with a similar scale in both understanding and generation tasks, with 7.1% higher Desc-GPT and 6.6% higher VQA-score, respectively. Codes and datasets are available at UniF2ace.
1 Introduction
Recently, unified multimodal models (UMMs) have emerged as a vibrant research area enabling cross-modality understanding and generation within a single “any-to-any” framework, marking a significant step toward artificial general intelligence (AGI) (Wu et al., 2024a; Shi et al., 2024; Li et al., 2024a; Zhou et al., 2024; Team, 2024; Xie et al., 2024). Given the central role of faces in daily life, applying this unified paradigm to achieve fine-grained face understanding and generation is essential for developing human-centric AGI. The practical applications are vast and critical: accurate face understanding is pivotal for identity verification (Srinivasan et al., 2024; Roshdy et al., 2024) and human-computer interaction (Liu, 2024; Chowdary et al., 2023), while high-fidelity face generation drives progress in creative industries (Melnik et al., 2024), virtual avatars (Yan et al., 2024), and data augmentation for model robustness (Melzi et al., 2023). These demanding real-world needs urge facial research to push the boundaries of multimodal understanding and generative modeling.
Figure 1: UniF2ace is the first unified multimodal model designed for face understanding and generation, encompassing tasks such as visual question answering(VQA) and text-to-image generation. The generated responses and images demonstrate UniF2ace’s potential in fine-grained face attributes.
As shown in Fig. 1, despite the critical importance of human faces, existing research faces two fundamental and intertwined challenges: First, the field remains fragmented, with current methodologies treating face understanding and generation as separate endeavors, failing to unify these capabilities into a single framework. Typically, face understanding models are often based on fine-tune pretrained multimodal large language models (MLLMs) on facial images with coarse text descriptions (Chettaoui et al., 2025; Sun et al., 2024a; Xing et al., 2024). Face generation models (Huang et al., 2023; Nair et al., 2023; Kim et al., 2024) often incorporate visual information, such as semantic masks and sketches, to guide high-fidelity face synthesis, but they cannot achieve direct generation from detailed captions to faces. This leads to disjointed workflows that are both computationally inefficient and functionally restrictive. Crucially, the absence of a unified framework represents a significant hurdle towards the realization of AGI within the domain of human faces. Second, there is a pervasive lack of fine-grained information processing across both understanding and generation tasks. This challenge stems from three problems: (a) The discrete diffusion model inherits the advantages of diffusion for image generation while enabling better scalability modeling with text tokens in UMMs (Yang et al., 2025; Xie et al., 2024). However, its specific implementation primarily relies on masked generative models (Chang et al., 2022), lacking a combination with accurate score matching (Lou et al., 2024), making it challenging to generate precise image details; (b) detailed attribute representations are prone to being discarded during the learning evolution in multimodal models (Zeng et al., 2024; He et al., 2025); and (c) the inaccessibility of cross-modality facial datasets featuring fine-grained attributes. Existing text-face datasets fall into two types: web-scraped low-resolution facial images with inaccurate captions (Li et al., 2024c; Zheng et al., 2022), and close-up facial datasets with limited attributes per caption (only 2 to 7) (Xia et al., 2021; Yu et al., 2023a), which lack detail. Moreover, current facial datasets do not include VQAs, limiting their use for fine-grained understanding tasks. Furthermore, this deficiency directly impacts high-quality face generation (Xiao et al., 2025; Deng et al., 2025; Wang et al., 2025).
To handle these challenges, we propose UniF2ace (see Fig. 1), the first UMM specifically tailored for unified and fine-grained face understanding and generation. UniF2ace aims to address the aforementioned critical challenges by simultaneously performing both tasks and capturing fine-grained facial attributes within a single model. Specifically, we firstly introduce a Dual Discrete Diffusion (D3Diff) loss within a novel theoretical proof that optimizes the negative log-likelihood, significantly improving generation quality. After that, we propose an integrated token-level and sequence-level Mixture-of-Experts(MoE) architecture that adaptively handling semantic and identity facial embeddings, effectively addressing the attribute forgotten phenomenon in representation evolvement and specialized fine-grained representation learning for both understanding and generation tasks. Finally, recognizing the critical role of data, we construct UniF2aceD-1M, a large-scale, specialized dataset containing 130K facial image-text pairs and 1M visual question-answering (VQA) pairs, with 17.7 attributes per caption. Extensive experiments on UniF2aceD-1M and other benchmarks demonstrate that UniF2ace significantly outperforms various top-leading single-task models or UMMs with a similar scale and dedicated face models across both understanding and generation tasks, with 7.1% higher Desc-GPT and 6.6% higher VQA-score. Besides, our method achieves comparable or even better accuracy than larger-scale models and establishes a strong baseline. Our main contributions are as follows:
  • A unified face understanding and generation framework: We introduce UniF2ace, the first unified multimodal model for fine-grained face understanding and generation, establishing a solid baseline.
  • A novel Dual Discrete Diffusion (D3Diff) loss function and a hybrid MoE architecture: We introduce D3Diff, a novel loss function within that theoretically unifies score-based diffusion and masked generative models, leading to a better approximation of the negative log-likelihood for high-fidelity generation and fine-grained attribute control. Additionally, we explore a hybrid Mixture-of-Experts (MoE) architecture at the token and sequence levels, adaptively incorporating the semantic and identity facial embeddings to complement the attribute forgotten phenomenon in representation evolvement.
  • We construct UniF2aceD-1M, a dataset containing 1M VQAs with an automated pipeline. Extensive experiments demonstrate that UniF2ace significantly outperforms or is on par with existing state-of-the-art models with a similar scale on various benchmarks, all while providing a more unified and efficient solution.
2 Related Work
The field of Unified Multimodal Models (UMMs) has seen significant progress in integrating diverse understanding and generation tasks within single frameworks for generic domains (Ma et al., 2024; Team, 2024; Xie et al., 2024). However, their application to fine-grained visual analysis, especially in the complex domain of human faces, remains largely unexplored. Within the face domain, existing research is primarily fragmented into separate understanding models (often MLLM-based) (Sun et al., 2024a; Xing et al., 2024) and generation models (typically diffusion-based) (Dai et al., 2025; Wang et al., 2024b). Crucially, these approaches often struggle with fine-grained attribute processing and fail to unify understanding and generation effectively. This dual deficiency represents a significant gap that UniF2ace addresses. We also provide a more comprehensive review of Unified Multimodal Models and Face Multimodal Models in Appendix  A.
3 METHODOLOGY
We introduce our unified multimodal model, UniF2ace, designed to model both the understanding and generation of fine-grained facial attributes. Our approach is realized from two perspectives: loss function (Sec. 3.1) and network architecture (Sec. 3.2). Regarding the generation strategy, we recognize that the generation of fine-grained facial attributes is significantly more challenging than understanding tasks, as highlighted in prior studies (Du et al., 2017; Zhou et al., 2024; Xie et al., 2024). To address this, we harness the theory of score matching in discrete diffusion (Lou et al., 2024) and propose the dual discrete diffusion (D3Diff) training strategy, ensuring the meticulous synthesis of facial details. For network architecture, existing UMMs typically focus on dense architectures (Zhou et al., 2024; Xie et al., 2024) or solely on achieving token-level MoE (Deng et al., 2025), lacking the selective integration of instance features. To overcome these limitations, we introduce token-level and sequence-level MoE layers. Distinct MoE modules are designed for generation and understanding tasks, selectively integrating information such as facial embeddings to enhance the model’s ability to capture subtle facial attributes.
3.1 Dual Discrete Diffusion
In generative modeling, masked generative models (Chang et al., 2022) are a widely adopted approach. However, in this section, we introduce discrete score matching and theoretically prove that it offers a better approximation to the negative log-likelihood. We also establish a theoretical connection between the two approaches and finally propose a new loss function to ensure stable optimization, thereby improving the alignment between the generated faces and fine-grained attributes in prompts.
In a discrete diffusion process, each image $\mathbf{x}_{0}$ is associated with a probability $q(\mathbf{x}_{0})$, and its latent distribution at time $t$ under noise adding is denoted by $q(\mathbf{x}_{t})$. The forward diffusion is modeled as a continuous-time Markov chain, governed by the linear ordinary differential equation (ODE):
$\frac{d}{dt}q_{t\mid s}({\bm{y}}\mid{\bm{x}})=q_{t\mid s}({\bm{y}}\mid{\bm{x}})\,{\bm{Q}}_{t},$ (1)
which converges to a stationary distribution as $t\rightarrow\infty$. Here, ${\bm{Q}}_{t}$ denotes a time-dependent sequence of transition matrices. The closed-form solution of this ODE can be expressed as ${\bm{Q}}_{t\mid s}=\mathbf{\exp}\big((\bar{\sigma}(t)-\bar{\sigma}(s))\,{\bm{Q}}\big)$, where $\bar{\sigma}(t)=\int_{0}^{t}\sigma(s)\,ds$ represents the cumulative noise level and $\mathbf{\exp}$ is the matrix exponential. The reverse process is given by Lou et al. (2024):
$\frac{dq_{T-t}}{dt}=\tilde{{\bm{Q}}}_{T-t}\,q_{T-t},\quad\tilde{{\bm{Q}}}=\frac{q_{t}({\bm{y}})}{q_{t}({\bm{x}})}\,{\bm{Q}}_{t}({\bm{x}},{\bm{y}})\,$, (2)
where $\tilde{{\bm{Q}}}$ is the reverse diffusion matrix. In our work, we focus on the absorbing state, which is widely used in masked generative models (Chang et al., 2022; Xie et al., 2024). Assuming independence among tokens, as supported by Sahoo et al. (2024); Shi et al. (2025), the exact formulation is deferred to Appendix D. The score-based discrete diffusion model (Lou et al., 2024) introduces a training-stable loss $\mathcal{L}_{\text{score}}(s_{\theta})$ to estimate the denoising score. It is defined as:
$\mathcal{L}_{\text{score}}(s_{\theta})=\mathbb{E}_{\mathbf{x}\sim p}\!\left[\!\sum_{\mathbf{y}\neq\mathbf{x}}w_{xy}\!\left(\!s_{\theta}(\mathbf{x})_{y}\!-\!\frac{q(\mathbf{y})}{q(\mathbf{x})}\log s_{\theta}(\mathbf{x})_{\mathbf{y}}\!+\!K\!\left(\frac{q(\mathbf{y})}{q(\mathbf{x})}\!\right)\right)\!\right],$ (3)
where $s_{\theta}({\mathbf{x}}_{t},t)\approx\left[\frac{q_{t}({\bm{y}}_{t})}{q_{t}({\bm{x}}_{t})}\right]_{{\bm{y}}_{t}\in\mathcal{X}}$ is the predicted score from the neural network, and $K(a)=a(\log a-1)$ is a normalizing constant ensuring $\mathcal{L}_{\text{score}}\geq 0$.
To illustrate the advantage of $\mathcal{L}_{\text{score}}$, we start from the negative log-likelihood (NLL), which serves as a fundamental criterion for evaluating the quality of training in generative models. Since exact computation of the NLL is generally infeasible, prior works have derived two different surrogate formulations that upper bound the NLL while remaining computationally tractable. Specifically, one is $\mathcal{L}_{1}=\mathcal{L}_{\text{score}}(s_{\theta})+D_{\text{KL}}\!\left(q_{T|0}(\cdot|\mathbf{x}_{0})\,\|\,p_{\text{base}}\right)$, which predicts the score (Lou et al., 2024). And the other is $\mathcal{L}_{2}=-\sum_{t=1}^{T}\mathbb{E}_{q(\mathbf{x}_{0})\,q(\mathbf{x}_{t}|\mathbf{x}_{0})}\!\big[\log p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{t})\big]-C$, widely used to predict masked tokens (Xie et al., 2024). where $C$ is a residual constant independent of the model parameters (see Appendix E for details). The following theorem formally establishes the relationship between these two surrogates and demonstrates that $\mathcal{L}_{1}$, which incorporates the score loss, yields a tighter upper bound on NLL.
Theorem 1.
Let $\log p_{\theta}({\mathbf{x}}_{0})$ denote the negative log-likelihood of the original data distribution. Then the following inequality holds:
$-\log p_{\theta}({\mathbf{x}}_{0})\leq\mathcal{L}_{1}\leq\mathcal{L}_{2}\,.$ (4)
The proof is deferred to Appendix E. Importantly, this result implies that $\mathcal{L}_{\text{score}}$ provides a tighter relaxation of the maximum likelihood objective compared to the masked generative related loss $\mathcal{L}_{2}$, thereby offering a more precise approximation of the NLL. In practice, the marginal distribution $q({\mathbf{y}})$ is often intractable, and the exact analytical form of $q({\mathbf{x}})$ is unknown. A key insight is that, in masked generative models, the posterior probability model $p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{t})$ can be related to the discrete diffusion score function via Bayes’ theorem:
$p_{\theta}({\mathbf{x}}_{0}\!\mid\!\mathbf{x}_{t})\approx q_{t}(\mathbf{x}_{t}\!\mid\!\mathbf{x}_{0})\!\left[\frac{q_{t}(\mathbf{x}_{0})}{q_{t}(\mathbf{x}_{t})}\!\right]_{\theta}\!\!\!=\!q_{t}(\mathbf{x}_{t}\!\mid\!\mathbf{x}_{0})\,\!s_{\theta}(\mathbf{x}_{t}).$ (5)
Leveraging this relation, we propose the dual discrete diffusion (D3Diff) loss for training posterior networks:
$\mathcal{L}_{\text{D3Diff}}=-\sum_{t=1}^{T}\mathbb{E}_{q(\mathbf{x}_{0})q(\mathbf{x}_{t}|\mathbf{x}_{0})}\left[\log p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{t})\right]+\alpha\mathcal{L}_{\text{score}}\left(p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{t})/q_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})\right),$ (6)
where $q(\mathbf{x}_{0})$ is the data distribution, $q(\mathbf{x}_{t}|\mathbf{x}_{0})$ is the forward diffusion distribution, and $p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{t})$ is the learned posterior parameterized by $\theta$. The score loss $\mathcal{L}_{\text{score}}$ is weighted by a hyperparameter $\alpha$. Eq. 6 establishes a computationally tractable connection between masked generative models and score-based models. Unlike traditional masked generative losses, which rely solely on likelihood, our D3Diff loss jointly optimizes two distinct upper bounds of the maximum likelihood objective, enabling stable optimization and fine-grained generation (See Tab. 5).
Figure 2: Our UniF2ace centered on two key innovations. First, we design the Transformer with Mixture-of-Experts (MoE) hierarchy: a token-level MoE provides task-specific routing for individual tokens, while a sequence-level MoE injects holistic, domain-specific features. Second, the model’s generative capability is optimized by our proposed D3Diff loss, which unifies masked generation with score matching to ensure high-fidelity synthesis of fine-grained facial details.
3.2 Multi-level Grouped Mixture-of-Expert
To capture fine-grained facial attributes while maintaining facial embeddings, we design distinct MoE layers, termed Multi-level Grouped MoE, tailored for both generation and understanding subtasks. This ensures optimal performance for each task, as illustrated in Fig. 2. We incorporate a sequence-level MoE layer after the token-level MoE layer to effectively process instance-level inputs, such as images and facial embeddings.
Token-Level MoE. We partition a feedforward neural network (FFN) into multiple experts with reduced hidden dimensions and use a Top-K activation strategy (Fig. 2). We also employ integrate generalized knowledge across contexts. Unlike prior methods, we introduce grouped MoE, dividing experts into two groups based on the different tasks of Text-to-Image (T2I) and Multimodal Understanding (MMU). Each group combines shared and routed MoE, with expert-level balance loss computed independently per group:
$\mathcal{L}_{\text{Balance}}=\lambda_{\text{t2i}}\sum_{i=1}^{N_{\text{t2i}}}f_{i}P_{i}+\lambda_{\text{mmu}}\sum_{j=1}^{N_{\text{mmu}}}f_{j}P_{j},$ (7)
where $\lambda_{\text{t2i}}$ and $\lambda_{\text{mmu}}$ are balance factors; $N_{\text{t2i}}$ and $N_{\text{mmu}}$ means the number of routed experts for T2I and MMU tasks, respectively; $f_{i}$ and $P_{j}$ denote expert selection frequency and probability.
Sequence-Level MoE. We propose sequence-level MoE, where distinct experts process the entire
Figure 3: Clip/Face Expert enhances the model’s understanding of fine-grained facial attributes by incorporating semantic and identity features.
image feature. We design three experts for the T2I group: copy expert (skip operation), zero expert (discard operation), and noise expert. The copy and zero experts require no additional parameters.
${\bm{E}}_{\text{copy}}({\bm{x}})=\mathbf{x}\quad\text{and}\quad{\bm{E}}_{\text{zero}}({\bm{x}})=\mathbf{0},$ (8)
where ${\bm{E}}_{\text{copy}}(\cdot)$ is the copy expert and ${\bm{E}}_{\text{zero}}(\cdot)$ is the zero expert. For the noise expert ${\bm{E}}_{\text{noise}}(\cdot)$, we first integrate the time-step embedding, which operates on the noise level $\bar{\sigma}(t)$ to obtain the noise embedding vector $\mathbf{v}_{\text{noise}}$, following score-based discrete diffusion models (Lou et al., 2023). Then, a resampler $\mathcal{S}:\mathbb{R}^{h}\rightarrow\mathbb{R}^{L\times D}$ maps ${\bm{v}}_{\text{noise}}$ into the sequence feature space (see Appendix E for resampler details). The resampled noise embedding is added as a matrix to the sequence feature. Formally, the noise expert’s output is:
${\bm{E}}_{\text{noise}}(\mathbf{x})=w\mathbf{x}+(1-w){\mathcal{S}({\bm{v}}_{\text{noise}})},$ (9)
$w=\text{Softmax}({\bm{W}}_{\text{noise}}\cdot\text{Flatten}(\mathbf{x})),$ (10)
where ${\bm{W}}_{\text{noise}}\in\mathbb{R}^{2\times(L\cdot D)}$ is a trainable weight matrix. In the MMU task, we include copy experts and introduce CLIP experts and face experts (See Figure 3), which are similar to noise experts. Next we extract image embeddings by CLIP (Radford et al., 2021) and face embeddings using AntelopeV2 as supplementary features to enhance fine-grained facial attribute capture. Formally, the outputs of the CLIP and face experts are:
${\bm{E}}_{\text{CLIP}}(\mathbf{x})=w_{clip}\mathbf{x}+(1-w_{clip})\mathcal{S}(\mathcal{G}({\bm{X}})),$ (11)
${\bm{E}}_{\text{face}}(\mathbf{x})=w_{face}\mathbf{x}+(1-w_{face})\mathcal{S}(\mathcal{F}({\bm{X}})),$ (12)
where $\mathcal{G}$ and $\mathcal{F}$ are the image encoder and face encoder, respectively. $\bm{X}$} is the input face image.
3.3 Overall Training Objectives
To perform both auto-regressive and discrete score-based diffusion modeling, we employ two learning objectives: 1) Next Token Prediction (NTP) and 2) Dual Discrete Diffusion. Given a sequence with $N$ image tokens $\mathcal{X}=\{{\bm{X}}_{1},{\bm{X}}_{2},\ldots,{\bm{X}}_{N}\}$ and $M$ text tokens $\mathcal{Y}=\{{\bm{Y}}_{1},{\bm{Y}}_{2},\ldots,{\bm{Y}}_{M}\}$. Then we maximize the likelihood of text tokens $\mathcal{Y}$ by employing the standard language modeling objective (NTP loss):
$\mathcal{L}_{\text{MMU}}=\sum_{i=1}^{M}\log P({\bm{Y}}_{i}\mid{\bm{Y}}_{<i},\mathcal{X}),$ (13)
Next, the overall training objectives of UniF2ace are formulated as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{MMU}}+\mathcal{L}_{\text{D3Diff}},$ (14)
4 Experiment
4.1 UniF2aceD-1M Dataset for Fine-grained Face understanding and generation
Dataset Face Resolution VQA Availability Image Caption Tokens (Avg.) Face Attributes (Avg. per Caption) LAION-Face (Zheng et al., 2022) Low 50M 16 2.7 FLIP-80M (Li et al., 2024c) 80M 22 4.4 FFHQ-Text (Zhou & Shimada, 2021) High 760 45 12.2 MM-CelebA-HQ (Karras et al., 2018) 30K 26 6.2 CelebV-Text (Yu et al., 2023b) 70K 80 4.3 UniF2aceD-1M (Ours) (1M) 130K 120 17.7
Figure 4: UniF2aceD-1M contains high-resolution facial images, the largest number of facial attributes, 130K fine-grained image-caption pairs and 1 million VQAs.
Existing datasets for multimodal facial modeling frequently suffer from significant limitations, hindering advancements in fine-grained understanding and generation. Common deficiencies include low-resolution imagery, imprecise or web-scraped captions lacking subtle attribute details, and a pervasive absence of comprehensive visual question-answering (VQA) pairs tailored to facial specifics (Li et al., 2024c; Zheng et al., 2022; Xia et al., 2021; Yu et al., 2023a; Karras, 2019). These shortcomings mean that current models struggle to synthesize nuanced facial expressions, comprehend intricate visual semantics, or reason effectively about complex facial attributes. To overcome these challenges and truly enable unified and fine-grained multimodal facial intelligence, we introduce UniF2aceD-1M. This high-quality dataset serves as a cornerstone of our framework, meticulously designed to bridge these critical data gaps.
As shown in Fig. 4, our UniF2aceD-1M provides a resource distinguished by its fine-grained detail and scale. It comprises nearly 130K high-fidelity facial images, each paired with richly detailed captions that encompass a wide spectrum of 46 attributes related to appearance, actions, and emotions. This meticulous level of detail is paramount for both robust model training and the generation of highly controllable and realistic facial outputs. Furthermore, a key innovation of UniF2aceD-1M is the inclusion of approximately 1M specialized VQA pairs. Unlike general VQAs, ours are meticulously crafted to probe diverse facial appearances, emotions, and provide detailed reasoning for character actions. This unique VQA collection is specifically designed to enhance MLLMs ability to understand and reason about fine-grained facial attributes through instruction tuning. By offering a substantial collection of high-quality facial imagery, richly detailed captions, and a unique, large-scale set of facial VQAs, UniF2aceD-1M sets a new standard, providing the indispensable resources for developing next-generation unified models capable of sophisticated fine-grained facial intelligence. More collection and operation details can be found in the Appendix B.
4.2 Metrics and Other Facial Datasets
We rigorously evaluate UniF2ace’s performance across both generation and understanding tasks on our UniF2aceD-1M test set. To provide a comprehensive assessment and verify the generalizability of our method, we also conduct evaluations on other public benchmarks, including FFHQ-Text (Zhou & Shimada, 2021), MM-CelebA (Xia et al., 2021), and CelebV-Text (Yu et al., 2023b). For generation tasks, we used VQAscore to measure the relevance of generated images to captions, reporting results based on CLIP-FlanT5-11B (VQAscore-CF5) (Lin et al., 2024b) and LLaVA-v1.5-13B (VQAscore-LV) (Liu et al., 2024c) for robust assessment. We also employ Fréchet Inception Distance (FID) to measure similarity to ground truth and VLM-score to evaluate facial realism. For understanding tasks, we follow LLaVA (Liu et al., 2023) and use GPT-4o (Hurst et al., 2024) and DeepSeek-v3 (Liu et al., 2024a) to score responses on a 1-10 scale across two dimensions: detailed captioning (Desc-GPT, Desc-DS), assessing accuracy in capturing face attributes, and VQA (Conv-GPT, Conv-DS), measuring precision in responding to fine-grained queries. To fully validate UniF2ace, we compare it with SOTA models. This includes generative models such as autoregressive LlamaGen (Yu et al., 2023c) and diffusion-based Stable Diffusion 3 (SD3) (Esser et al., 2024), as well as leading unified multimodal models (UMMs) like TokenFlow (Qu et al., 2024) and OmniFlow (Li et al., 2024a). More implementations details are in Appendix C.
4.3 Face Generation
Table 1: Comparison of face generation of UniF2ace with generative-only and UMMs. Bold indicates the best, while underlined denotes the best. We use red to highlight the larger-scale model.
[…]
Gen. Only LlamaGen (Sun et al., 2024b) AR 0.8B 0.746 0.551 183.466 49.773
DALL-E 3 (Betker et al., 2023) AR - 0.845 0.644 106.477 50.122
SD3 (Esser et al., 2024) Diff 2B 0.903 0.671 93.471 75.944
SDXL (Podell et al., 2023) Diff 2.6B 0.876 0.660 123.095 72.764
Flux.1-dev (Labs, 2024) Diff 12B 0.893 0.674 76.427 84.513
Und. and Gen. TokenFlow (Qu et al., 2024) AR 7B 0.871 0.664 98.194 73.177
OmniFlow (Li et al., 2024a) Diff 3.4B 0.798 0.585 180.933 24.960
JanusFlow (Ma et al., 2024) AR + Diff 1.3B 0.881 0.653 72.825 61.593
Show-o (Xie et al., 2024) AR + Diff 1.3B 0.855 0.650 142.557 75.618
UniF2ace(Ours) AR + Diff 1.8B 0.894 0.679 66.005 88.049
Table 2: Comparison of face generation on other public datasets. The experimental setup utilized the built-in short captions of datasets as text prompts for generation.
Type Model Params FFHQ-Text $M$-CelebA CelebV-Text
[…]
Gen.Only LlamaGen (Sun et al., 2024b) 0.8B 0.336 201.341 46.412 0.358 187.311 48.121 0.721 289.841 46.906
DALL-E 3 (Betker et al., 2023) - 0.385 196.132 49.131 0.413 158.795 49.130 0.792 295.131 54.359
SD3 (Esser et al., 2024) 2B 0.423 156.129 74.492 0.459 105.141 80.142 0.803 239.313 74.127
SDXL (Podell et al., 2023) 2.6B 0.396 181.261 64.255 0.420 139.028 73.149 0.788 271.319 70.991
Flux.1-dev (Labs, 2024) 12B 0.434 136.360 83.621 0.467 128.462 87.764 0.806 254.043 84.901
Gen.&Und. TokenFlow (Qu et al., 2024) 7B 0.409 160.023 74.349 0.421 129.562 71.092 0.781 273.972 79.526
OmniFlow (Li et al., 2024a) 3.4B 0.376 228.094 25.431 0.368 201.413 27.892 0.800 290.131 36.839
JanusFlow (Ma et al., 2024) 1.3B 0.413 149.231 60.984 0.445 129.131 63.418 0.797 259.236 66.587
Show-o (Xie et al., 2024) 1.3B 0.391 177.053 73.141 0.428 141.311 74.242 0.785 260.210 70.482
UniF2ace(Ours) 1.8B 0.451 125.287 87.412 0.481 85.179 86.978 0.804 224.412 94.986
Figure 5: Comparative analysis of face images generation quality across SDXL  (Podell et al., 2023), TokenFlow (Qu et al., 2024), OmniFlow  (Li et al., 2024a), Show-o (Xie et al., 2024), and UniF2ace. Our proposed UniF2ace effectively captures more detailed information from prompts.
Figure 6: Activation frequency of Token-Level and Sequence-Level MoE in different layers. The left column indicates understanding tasks, while the right column indicates generation tasks. Larger circles indicate experts who are activated more frequently.
Generation Performance on UniF2aceD-1M and Public Dataset. On our UniFaceD-1M benchmark (Tab. 1), our 1.8B parameter UniF2ace sets a new state-of-the-art, outperforming all competing UMMs on key generation metrics including FID, VQA-score, and VLM-score. Furthermore, the model also demonstrates robust generalization, consistently achieving leading scores on public cross-facial datasets such as FFHQ-Text (Zhou & Shimada, 2021), MM-CelebA (Xia et al., 2021), and CelebV-Text (Yu et al., 2023a) (Tab. 2). This strong and consistent performance validates the effectiveness of our D3Diff loss and multi-level grouped MoE architecture for high-quality, fine-grained facial image generation across diverse settings.
Visualization Analysis. As shown in Fig. 5, we conduct qualitative evaluation on challenging UniF2aceD-1M test scenarios involving complex facial details. UniF2ace excels at generating realistic faces that capture fine-grained details from complex prompts (e.g., ”rosy cheeks,” ”hoop earrings”), visibly outperforming other models. More examples can be found in Fig. 8 and Fig. 9. Besides, as shown in Fig. 6, we analyze MoE activation frequencies across layers. For token-level MoEs, high activation frequencies are concentrated between experts 5 and 8, indicating limited token feature variability in the generation task. For sequence-level MoEs, noise and zero expert activations are evenly distributed, indicating effective training with selective noise embedding and truncation.
4.4 Face Understanding
Table 3: Comparison of face understanding of UniF2ace with understanding-only and UMMs.
[…]
Und. Only VILA1.5 (Lin et al., 2023) AR 3B 4.76 5.20 6.56 6.54
Qwen2-VL (Wang et al., 2024a) AR 7B 5.16 6.27 5.50 6.86
LLaVA-v1.5 (Liu et al., 2024b) AR 7B 4.28 5.48 4.84 6.20
InternVL2.5 (Chen et al., 2024) AR 8B 5.62 5.89 6.30 6.55
Qwen2.5-VL (Bai et al., 2025) AR 3B 4.88 6.38 4.98 6.75
Und. and Gen. TokenFlow (Qu et al., 2024) AR 7B 5.02 5.80 5.82 6.39
OmniFlow (Li et al., 2024a) Diff 3.4B 1.62 - 1.90 -
JanusFlow (Ma et al., 2024) AR + Diff 1.3B 4.88 6.06 5.42 6.77
Show-o (Xie et al., 2024) AR + Diff 1.3B 3.88 4.17 5.24 4.90
UniF2ace(Ours) AR + Diff 1.8B 6.02 6.53 7.38 7.29
Table 4: Comparison of face understanding on other public datasets. The experiments utilized the dataset’s captions as labels for captioning task evaluation, showing the robustness of UniF2ace.
Type Model Params FFHQ-Text $M$-CelebA CelebV-Text
Desc-GPT Desc-DS Desc-GPT Desc-DS Desc-GPT Desc-DS
Und.Only VILA1.5 (Lin et al., 2023) 3B 4.29 4.79 4.48 4.59 4.61 4.76
Qwen2-VL (Wang et al., 2024a) 7B 4.68 5.41 5.11 5.40 4.90 4.95
LLaVA-v1.5 (Liu et al., 2024b) 7B 4.01 4.60 4.29 4.26 4.54 4.50
InternVL2.5 (Chen et al., 2024) 8B 5.09 5.58 4.75 4.98 5.07 5.01
Qwen2.5-VL (Bai et al., 2025) 3B 4.38 4.92 4.72 4.70 5.20 5.10
Gen.&Und. TokenFlow (Qu et al., 2024) 7B 5.04 5.75 4.99 5.01 4.86 5.10
OmniFlow (Li et al., 2024a) 3.4B 2.83 3.06 3.41 3.38 2.90 3.03
JanusFlow (Ma et al., 2024) 1.3B 4.31 5.15 4.60 4.71 4.54 4.86
Show-o (Xie et al., 2024) 1.3B 3.86 4.67 4.38 4.39 4.49 4.57
UniF2ace(Ours) 1.8B 5.12 5.92 6.24 6.80 5.87 5.29
Understanding Performance on UniF2aceD-1M and Public Dataset. On our UniF2aceD-1M benchmark (Tab. 3), UniF2ace sets a new state-of-the-art in fine-grained facial understanding, achieving the highest scores across all metrics. Crucially, it surpasses even larger, specialized models like InternVL2.5 (8B) and all competing UMMs. Furthermore, as shown in Tab. 4, this superior understanding capability demonstrates strong generalization, as UniF2ace also consistently achieves top captioning scores on public cross-facial datasets. This robust performance across diverse benchmarks validates the effectiveness of our approach in learning transferable, fine-grained facial representations, affirming UniF2ace’s leading position in comprehensive multimodal facial understanding.
Quantitative Evaluation. We compare UniF2ace with advanced autoregressive multimodal models like Qwen2-VL (Wang et al., 2024a) (7B) and similar UMMs (Wu et al., 2024c; Xie et al., 2024).
Visual comparisons (right Fig.) confirm UniF2ace’s superior fine-grained understanding in VQA tasks, where it accurately identifies subtle details like ”sideburns” and ”earrings” that competitors miss. We also provide more examples for captioning in the Appendix (Fig. 10). Besides, as shown in Fig. 6 (left column), we analyze MoE activation frequencies in the understanding task. Token-level MoEs often select the same expert in the top (the closest to the prediction head) and bottom layers. For sequence-level MoEs, face and CLIP experts are more frequently activated in layers closer to the top, indicating that deeper layers benefit from visual embeddings to better understand face images. Notably, activation patterns are different across the generation groups, highlighting the effectiveness of our group-based strategy.
4.5 Ablation Studies
[…]
Table 5: Ablation study with different loss weights.
[…]
score-matching and masked generative losses. The complete D3Diff loss significantly outperforms using either loss component individually. Crucially, the superiority of the score-only loss over the masked-only loss empirically validates our theoretical analysis in Appendix E.
Table 6: Ablation of Face and CLIP Expert.
[…]
Table 7: Ablation of Top-k in Seq-level MoE.
[…]
Ablation of MoEs Architecture. We ablate our multi-level MoE design ( Tab. 4.5, Tab. 7, Tab. 8). The results (Tab. 8) first confirm that combining token-level and sequence-level MoEs achieves the best
Table 8: Ablation study of token- and sequence-level MoE.
[…]
performance, with each component individually outperforming the non-MoE baseline. We further analyze the sequence-level MoE, finding that: (1) for understanding tasks, using both CLIP and Face experts is optimal in fine-grained facial understanding (Tab. 4.5); and (2) a Top-k=2 selection strategy provides the best balance between performance and efficiency (Tab. 7). These findings validate our hierarchical and specialized MoE design.
5 Conclusion
This paper introduces UniF2ace, the first unified multimodal model (UMM) designed for fine-grained face understanding and generation. The model bridges the gap between score-based models and masked generative models in discrete diffusion, while leveraging token-level and sequence-level mixture-of-experts (MoE) to sparsify the model. Extensive experiments show that UniF2ace outperforms existing UMMs and even surpasses larger generation-only or understanding-only models. This underscores the potential of our improvements to guide future research in face applications of UMM. Additionally, we constructed a large-scale face-text aligned dataset, UniF2aceD-1M, to further advance multimodal research in the community.
Appendix A Related Works
Unified Multimodal Models.
Recent works (Ma et al., 2024; Li et al., 2024d; Wu et al., 2024a; Chen et al., 2025; Wang et al., 2024c) in image understanding and generation have primarily focused on unified multimodal models (UMMs). Early approaches (Li et al., 2024b; Wu et al., 2024b) often integrated external decoders of diffusion models (DMs) with text autoregressive models (ARMs). Inspired by next-token prediction tasks, they proposed using a single Transformer (Vaswani et al., 2017) model to unify understanding and generation (Wu et al., 2024c). For instance, Janus-Pro (Chen et al., 2025) decouples the visual encoder into specialized tokenizers for separate handling of understanding and generation tasks. Chameleon (Team, 2024) and Emu3 (Wang et al., 2024c) employ an ARM to simultaneously manage both tasks, highlighting the advantages of autoregressive models in multitask settings. Additionally, Transfusion (Zhou et al., 2024) and Show-o (Xie et al., 2024) combine a text ARM with a visual DM, enabling seamless integration of image understanding and generation. These studies have advanced the fusion of visual and text generation models, enhancing performance on multimodal tasks. However, despite the proliferation of UMMs, their application has largely been limited to generic domain tasks, with limited exploration in fine-grained visual analysis, particularly in the face domain. Unlike previous UMMs that simply combine ARMs and DMs, we pioneer sparse UMMs by introducing both token-level and sequence-level Mixture of Experts (MoEs), significantly improving model performance.
Face Multimodal Models.
Face multimodal models are primarily categorized into two types: face understanding models and face generation models. For understanding, early models were task-specific and lacked multimodality (Miyato et al., 2018; Zhang et al., 2024; Wang et al., 2023; Lee et al., 2023). Recent works (Chettaoui et al., 2025; Sun et al., 2024a; Xing et al., 2024; Zhao et al., 2024) leverage the reasoning capabilities of LLMs or MLLMs, often using MLLM-generated face Q&A data to fine-tune or post-train foundation models, incorporating face domain knowledge. For example, EMO-LLaMA (Xing et al., 2024) introduces facial experts to extract facial features, which are aggregated with handcrafted prompts and fed into LLaMA (Touvron et al., 2023), enabling it to answer facial-related queries. For generation, recent works (Dai et al., 2025; Wang et al., 2024b; Huang et al., 2023; Kim et al., 2024) focus on using diffusion models to personalize face images by conditioning on textual and visual information, such as semantic masks, but avoid directly capturing fine-grained face attributes from text prompts. Despite these advances in understanding and generation separately, developing unified multimodal models (UMMs) remains a significant research challenge. Addressing this gap can enhance cross-modal capabilities and advance progress toward Artificial General Intelligence (AGI).
Appendix B Dataset Construction
To overcome the limitations of existing datasets in the realm of multimodal facial modeling, we introduce a high-quality dataset called UniF2aceD-1M, which boasts a remarkable alignment between facial images and textual descriptions (see Fig. 7). This dataset encompasses nearly 130K facial images, each paired with richly detailed captions. Additionally, it contains approximately 1M visual question answers, significantly enhancing its value for training and evaluating multimodal models. By offering such a comprehensive resource, we aim to propel advancements in facial image understanding and generation, establishing a solid foundation for a wide range of multimodal learning tasks. The creation of UniF2aceD-1M encompassed three key stages. (1) Step-1: Collect high-quality facial images. (2) Step-2: Generate detailed captions. (3) Step-3: Create question-answering pairs. Each stage is outlined in detail below.
(1) Step-1: Collect High-quality Facial Images.
In this step, we curated more than 130,000 high-quality facial images from the following distinguished datasets. CelebV-HQ (Zhu et al., 2022) is a large-scale video dataset featuring 35,666 clips representing 15,653 identities, each clip meticulously annotated with 83 facial attributes. We extracted one key frames from each video to utilize detailed annotations for fine-grained face-text alignment. Flickr-Faces-HQ (FFHQ) (Karras, 2019) provided 70,000 high-quality PNG images at a resolution of 1024 by 1024, offering substantial diversity in attributes such as age and ethnicity. Multi-Modal-CelebA-HQ (MM-CelebA-HQ) (Xia et al., 2021) contributed 30,000 high-resolution images paired with descriptive captions that have proven invaluable for facial generation and analysis.
Figure 7: Pipeline and examples of UniF2aceD-1M construction. Left: A three-stage pipeline for building UniF2aceD-1M. Step-1: High-quality face images are collected. Step-2: Detailed captions are generated by GPT-4o with a face attribute model trained to classify fine-grained appearance, action, and emotion. Step-3: Question-answering pairs are created. These stages collectively refine GPT-4o-generated captions and produce fine-grained descriptions for VQAs generation. Right: A representative example showcasing UniF2aceD-1M’s ability to correct (e.g., gender), enhance (e.g., bags under eyes), and reason (e.g., talking, slight tiredness) in GPT-4o-generated captions.
(2) Step-2: Generate Detailed Captions.
Existing face image datasets often lack detailed descriptions of fine-grained attributes like bags under eyes or jewelry. To handle this, we develop a two-stage caption generation process.
In Stage I, we employed an advanced MLLM such as GPT-4o (Hurst et al., 2024) to produce initial captions. We designed a specialized prompt that incorporated brief face descriptions from the MM-CelebA-HQ dataset (Xia et al., 2021) to help GPT-4o accurately describe key facial attributes, including appearance, emotion, and actions. The detailed descriptions of all prompts are presented later (see Fig. 11).
In Stage II, we refined these captions by training face attribute classification models using the CelebV-HQ dataset (Zhu et al., 2022). Focusing on single-person images, we used the pretrained face model AntelopeV2111https://github.com/deepinsight/insightface to extract face embeddings. By combining these with image embeddings from CLIP (Radford et al., 2021), we trained classification heads for appearance, action, and emotion attributes. We selected 29 appearances with accuracies over 93%, 10 actions with accuracies over 87%, and 7 emotions with accuracies over 80% as final preditions for inference. These highly accurate attributes were then predicted for all remaining images in FFHQ and MM-CelebA-HQ datasets (Karras, 2019; Xia et al., 2021). Finally, a prompt integrating these classification results with the Stage I captions was fed into GPT-4o to generate final captions that are both highly accurate and diverse.
(3) Step-3: Create Question-answering Pairs.
In this step, we proposed 1M VQAs covering diverse facial appearances, emotions, and character action reasoning for our UniF2aceD-1M dataset. These VQAs are designed to enhance MLLMs’ ability to understand fine-grained facial attributes through instruction tuning. Inspired by LLaVA (Liu et al., 2024c), we carefully designed prompts to enable GPT-4 (Achiam et al., 2023) to generate a series of VQAs based on image captions, facilitating fine-grained understanding and reasoning. Most current face-text datasets lack VQAs, while VQAs in general image-text datasets often focus on people’s clothing, location, and behavior, neglecting detailed facial descriptions. In contrast, our proposed VQAs encompass diverse facial details, including hair, nose, eyes, mouth, ears, skin, eyebrows, and adornments. Additionally, since facial attributes can reflect a character’s ongoing actions, our VQAs incorporate detailed reasoning processes to infer and describe these actions. By organizing the VQAs into the same format as the LLaVA dataset (Liu et al., 2024c), we streamlined the process of adapting multimodal face models for post-training. This alignment minimizes alteration costs, ensuring efficient integration and enabling the models to leverage both datasets seamlessly for improved performance.
Appendix C Implementations Details
We train our model on the UniF2aceD-1M training dataset part, comprising 120K 256 $\times$ 256 face images, each annotated with detailed captions and seven to eight VQAs, about 900K. UniF2ace utilizes discrete image tokens as input, represented by the pre-trained MAGVIT-v2 (Yu et al., 2023c). For token-level MoE, each group (generation and understanding tasks) includes one shared expert and eight routed experts, selected via a top-2 strategy. The expert structure is a single-layer MLP with the gating mechanism (Dai et al., 2024). In sequence-level MoE, the generation group employs two copy experts, one zero expert, and one noise expert. Noise embedding is implemented using sinusoidal embedding, following  (Nichol et al., 2021). The noise resampler uses a 4-layer Multi-Head Attention mechanism to map noise embeddings to the UniF2ace hidden space. For the understanding group, there are two copy experts, one CLIP expert, and one face expert. We use CLIP-ViT for image embedding and AntelopeV2 for face embedding, with the resampler configuration matching that of the noise expert. Moreover, training is divided into two stages: Stage I uses only captions for generation and understanding tasks, while Stage II incorporates VQAs into the understanding task. This pipeline transitions the model from general image feature understanding to fine-grained feature capture. Both stages are trained on 8 NVIDIA A100 (80GB) GPUs, optimized using AdamW with a weight decay of 0.01, 5K warm-up steps, and an initial learning rate of 5e-5 with cosine scheduling. The total batch size is 600 for Stage I and 480 for Stage II, with 20K steps for Stage I and 40K steps for Stage II. In the inference process of UniF2ace, following the computation method in  (Lin et al., 2024a), we compute the maximum and minimum activation parameters for UniF2ace under the Top-2 strategy due to the different number of parameters included between different experts in the sequence-level MoE. The total number of parameters for UniF2ace is 1.84B, the maximum activation parameter is about 1.63B, and the minimum activation parameter is about 1.42B. The average number of activation parameters tested in the UniF2aceD-1M test dataset is 1.47B.
Appendix D Absorbing-state Case with Independence between Tokens.
The absorbing-state case means that for any single token $x$ with possible values in $\mathcal{X}=$ $\{1,\ldots,N\}$, the transition matrix is
$\bm{Q}^{\text{absorb }}=\left[\begin{array}[]{ccccc}-1&0&\cdots&0&1\\ 0&-1&\cdots&0&1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&-1&1\\ 0&0&\cdots&0&0\end{array}\right]\,.$ (20)
The reverse transition rate matrix of the reverse process from state $\mathbf{x}_{t}$ to state $\mathbf{\hat{x}}_{t}$ is
$\tilde{\bm{Q}}_{t}\left(\mathbf{x}_{t},\mathbf{\hat{x}}_{t}\right)=\begin{cases}\frac{q_{t}\left(\mathbf{\hat{x}}_{t}\right)}{q_{t}\left(\mathbf{x}_{t}\right)}\bm{Q}_{t}\left(\mathbf{\hat{x}}_{t},\mathbf{x}_{t}\right),&\mathbf{\hat{x}}_{t}\neq\mathbf{x}_{t}\\ -\sum_{k\neq\mathbf{x}_{t}}\tilde{\bm{Q}}_{t}\left(\mathbf{x}_{t},k\right),&\mathbf{\hat{x}}_{t}=\mathbf{x}_{t}\end{cases}\,.$ (21)
As $\bm{Q}_{t}\left(\mathbf{\hat{x}}_{t},\mathbf{x}_{t}\right)$ is known, it is sufficient to estimate the concrete score $\frac{q_{t}\left(\mathbf{\hat{x}}_{t}\right)}{q_{t}\left(\mathbf{x}_{t}\right)}$ by a score network $s_{\theta}\left(\mathbf{x}_{t},t\right)\approx$ $\left[\frac{q_{t}\left(\mathbf{\hat{x}}_{t}\right)}{q_{t}\left(\mathbf{x}_{t}\right)}\right]_{\mathbf{\hat{x}}_{t}\in\mathcal{X}}$. Score based discrete diffusion model is an effective objective to train the score network (Meng et al., 2022; Lou et al., 2023). Specifically, the score function in a multidimensional discrete space is
$\bm{s}_{\theta}\left(\bm{x}_{t},t\right)_{\hat{\bm{x}}_{t}}=\bm{s}_{\theta}\left(\mathbf{x}_{t}^{1}\ldots\mathbf{x}_{t}^{i}\ldots\mathbf{x}_{t}^{d},t\right)\left[i,\widehat{\mathbf{x}}_{t}^{i}\right]\approx\frac{q_{t}\left(\mathbf{x}_{t}^{1}\ldots\widehat{x}_{t}^{i}\ldots\mathbf{x}_{t}^{d}\right)}{q_{t}\left(\mathbf{x}_{t}^{1}\ldots\mathbf{x}_{t}^{i}\ldots\mathbf{x}_{t}^{d}\right)}\,$, (22)
and accordingly,
$\tilde{\bm{Q}}_{t}\left(\mathbf{x}_{t}^{1}\ldots\mathbf{x}_{t}^{i}\ldots\mathbf{x}_{t}^{d},\mathbf{x}_{t}^{1}\ldots\widehat{\mathbf{x}}_{t}^{i}\ldots\mathbf{x}_{t}^{d}\right)\approx\bm{Q}_{t}\left(\widehat{\mathbf{x}}_{t}^{i},\mathbf{x}_{t}^{i}\right)\bm{s}_{\theta}\left(\mathbf{x}_{t}^{1}\ldots\mathbf{x}_{t}^{i}\ldots\mathbf{x}_{t}^{d},t\right)\left[i,\widehat{x}_{t}^{i}\right]\,.$ (23)
Appendix E Proof of Theorem 1
To prove Theorem 1, we first introduce two loss formulas and establish useful lemmas.
  • (1)
    $\mathcal{L}_{1}=\mathcal{L}_{\text{score }}\left(s_{\theta}\right)+D_{KL}\left(q_{T\mid 0}\left(\cdot\mid\mathbf{x}_{0}\right)\|q_{\text{base }}\right)$, where $\mathcal{L}_{\text{score }}\left(\mathbf{x}_{0}\right)$ is the diffusion weighted denoising score entropy for data point $\mathbf{x}_{0}$, and $s_{\theta}=\frac{q_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{t})}{q(\mathbf{x}_{t}|\mathbf{x}_{0})}$
    $\mathcal{L}_{\text{score }}\left(s_{\theta}\right)=$ $\int_{0}^{T}\mathbb{E}_{\mathbf{x}_{t}\sim q_{t\mid 0}\left(\cdot\mid\mathbf{x}_{0}\right)}\sum_{\mathbf{y}\neq\mathbf{x}_{t}}Q_{t}\left(\mathbf{x}_{t},y\right)\left(s_{\theta}\left(\mathbf{x}_{t},t\right)_{\mathbf{y}}\right$. (24)
    $-\left.\frac{q_{t\mid 0}\left(\mathbf{y}\mid\mathbf{x}_{0}\right)}{q_{t\mid 0}\left(\mathbf{x}_{t}\mid\mathbf{x}_{0}\right)}\log s_{\theta}\left(\mathbf{x}_{t},t\right)_{\mathbf{y}}+K\left(\frac{q_{t\mid 0}\left(\mathbf{y}\mid\mathbf{x}_{0}\right)}{q_{t\mid 0}\left(\mathbf{x}_{t}\mid\mathbf{x}_{0}\right)}\right)\right)dt\,.$
  • (2)
    $\mathcal{L}_{2}=-\sum_{t=1}^{T}\mathbb{E}_{q\left(\mathbf{x}_{0}\right)q\left(\mathbf{x}_{t}\mid\mathbf{x}_{0}\right)}\left[\log p_{\theta}\left(\mathbf{x}_{0}\mid\mathbf{x}_{t}\right)\right]-C\,$, where $C$ is a constant independent of the model parameters. By (Xie et al., 2024), $C=C_{1}+C_{2}$, and constants $C_{1}$ and $C_{2}$ are shown as:
    $C_{1}$ $=\mathbb{E}_{q\left(\mathbf{x}_{0:T}\right)}\left[-\sum_{t=1}^{T}\log q\left(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}\right)+\underbrace{\log p\left(\mathbf{x}_{T}\right)}_{\text{Note that }p\left(\mathbf{x}_{T}\right)=q\left(\mathbf{x}_{T}\right)}\right]$ (25)
    $=\mathbb{E}_{q\left(\mathbf{x}_{0:T}\right)}\left[-\sum_{t=1}^{T}\log q\left(\mathbf{x}_{t},\mathbf{x}_{t-1}\right)+\sum_{t=0}^{T}\log q\left(\mathbf{x}_{t}\right)\right]$
    $C_{2}$ $=\mathbb{E}_{q\left(\mathbf{x}_{0:T}\right)}\left[\sum_{t=1}^{T}\log q\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}\right)\right]-\mathbb{E}_{q\left(\mathbf{x}_{0:T}\right)}\left[\sum_{t=1}^{T}\sum_{\tilde{\mathbf{x}}_{0}}q\left(\tilde{\mathbf{x}}_{0}\mid\mathbf{x}_{t-1}\right)\log q\left(\tilde{\mathbf{x}}_{0}\mid\mathbf{x}_{t}\right)\right]$
    $=\mathbb{E}_{q\left(\mathbf{x}_{0:T}\right)}\left[\sum_{t=1}^{T}\log q\left(\mathbf{x}_{t},\mathbf{x}_{t-1}\right)-\sum_{t=1}^{T}\log q\left(\mathbf{x}_{t}\right)\right]$
    $\quad\quad-\sum_{t=1}^{T}\mathbb{E}_{q\left(\mathbf{x}_{0:T}\right)q\left(\tilde{\mathbf{x}}_{0}\mid\mathbf{x}_{t-1}\right)}\left[\log q\left(\tilde{\mathbf{x}}_{0}\mid\mathbf{x}_{t}\right)\right]\,.$
    $C_{1}+C_{2}=\mathbb{E}_{q\left(\mathbf{x}_{0:T}\right)}\left[\log q\left(\mathbf{x}_{0}\right)-\sum_{t=1}^{T}\log q\left(\mathbf{x}_{0}\mid\mathbf{x}_{t}\right)\right].$ (26)
Let $L=\log p_{\theta}(\mathbf{x}_{0})$ be the model’s log-likelihood for a data point $\mathbf{x}_{0}$, and let $K$ be its variational lower bound:
$K=\mathbb{E}_{q\left(\mathbf{x}_{0}\right)q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}\left[\log\frac{p_{\theta}\left(\mathbf{x}_{0:T-1}\mid\mathbf{x}_{T}\right)}{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}+\log p\left(\mathbf{x}_{T}\right)\right]\,.$ (27)
Then, the following inequality chain holds:
$L\geq K=-\mathcal{L}_{1}\geq-\mathcal{L}_{2}\,$, (28)
where $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are defined as above.
The proof is based on two applications of Jensen’s inequality to the log-likelihood.
1. Proving $L\geq K$
This is the standard variational lower bound for diffusion models, derived by applying Jensen’s inequality:
$\log p_{\theta}(\mathbf{x}_{0})$ $=\log\int p_{\theta}(\mathbf{x}_{0:T})d\mathbf{x}_{1:T}$ (29)
$=\mathbb{E}_{q(\mathbf{x}_{0})}\left[\log\mathbb{E}_{q(\mathbf{x}_{1:T}\mid\mathbf{x}_{0})}\left[\frac{p_{\theta}(\mathbf{x}_{0:T-1}\mid\mathbf{x}_{T})p_{\theta}(\mathbf{x}_{T})}{q(\mathbf{x}_{1:T}\mid\mathbf{x}_{0})}\right]\right]$
$\stackrel{{\scriptstyle(a)}}{{\geq}}\mathbb{E}_{q(\mathbf{x}_{0:T})}\left[\log\frac{p_{\theta}(\mathbf{x}_{0:T-1}\mid\mathbf{x}_{T})}{q(\mathbf{x}_{1:T}\mid\mathbf{x}_{0})}+\log p_{\theta}(\mathbf{x}_{T})\right]=K\$
Here, we assume $p_{\theta}(\mathbf{x}_{T})\approx q(\mathbf{x}_{T})$. This proves the first part of the inequality.
2. Proving $K=-\mathcal{L}_{1}$
$K$ $=\mathbb{E}_{q(\mathbf{x}_{0:T})}\left[\log\prod_{t=1}^{T}\frac{p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})}{q(\mathbf{x}_{t}|\mathbf{x}_{t-1})}+\log p_{\theta}(\mathbf{x}_{T})\right]$ (30)
$=\sum_{t=1}^{T}\mathbb{E}_{q(\mathbf{x}_{0:T})}\left[\log\frac{p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})}{q(\mathbf{x}_{t}|\mathbf{x}_{t-1})}\right]+\mathbb{E}_{q(\mathbf{x}_{0:T})}\left[\log p_{\theta}(\mathbf{x}_{T})\right]$
From the derivations in (Sohl-Dickstein et al., 2015), the variational lower bound $K$ can be strictly rewritten in terms of KL divergences, which directly correspond to our $\mathcal{L}_{1}$. Specifically,
$K$ $=-\sum_{t=2}^{T}\int d\mathbf{x}_{0}d\mathbf{x}_{T}q\left(\mathbf{x}_{0},\mathbf{x}_{T}\right)\cdot\text{KL}\left(q\left(\mathbf{x}_{T-1}\mid\mathbf{x}_{T},\mathbf{x}_{0}\right)\|p\left(\mathbf{x}_{T-1}\mid\mathbf{x}_{T}\right)\right)$ (31)
$\quad\quad+H_{q}\left(\mathbf{x}_{T}\mid\mathbf{x}_{0}\right)-H_{q}\left(\mathbf{x}_{1}\mid\mathbf{x}_{0}\right)-H_{p}\left(\mathbf{x}_{T}\right)$
Since
$H_{q}\left(\mathbf{x}_{T}\mid\mathbf{x}_{0}\right)-H_{p}\left(\mathbf{x}_{T}\right)$ $=\int_{\mathbf{x}_{T}}\int_{\mathbf{x}_{0}}q\left(\mathbf{x}_{T}\mid\mathbf{x}_{0}\right)q\left(\mathbf{x}_{0}\right)\log q\left(\mathbf{x}_{T}\mid\mathbf{x}_{0}\right)d\mathbf{x}_{0}d\mathbf{x}_{T}$ (32)
$-\int_{\mathbf{x}_{T}}\int_{\mathbf{x}_{0}}q\left(\mathbf{x}_{t}\mid\mathbf{x}_{0}\right)q\left(\mathbf{x}_{0}\right)d\mathbf{x}_{0}\log p\left(\mathbf{x}_{T}\right)d\mathbf{x}_{T}$
$=\int_{\mathbf{x}_{T}}\int_{\mathbf{x}_{0}}q\left(\mathbf{x}_{T}\mid\mathbf{x}_{0}\right)q\left(\mathbf{x}_{0}\right)\log\frac{q\left(\mathbf{x}_{T}\mid\mathbf{x}_{0}\right)}{p\left(\mathbf{x}_{T}\right)}d\mathbf{x}_{0}d\mathbf{x}_{T}$
$=\mathbb{E}_{p_{\text{data}}(\mathbf{x}_{0})}\left[\text{KL}\left(q\left(\mathbf{x}_{T}\mid\mathbf{x}_{0}\right)\|p\left(\mathbf{x}_{T}\right)\right]\right$.\,,
and $H_{q}\left(\mathbf{x}_{1}\mid\mathbf{x}_{0}\right)=\mathbb{E}_{p_{\text{data }}}\mathbb{E}_{q\left(x_{1}\mid\mathbf{x}_{0}\right)}\left[\log p_{0\mid 1}\left(\mathbf{x}_{0}\mid x_{1}\right)\right]$, then the above formula is equivalent to
$-\mathbb{E}_{\mathbf{x}_{t}\sim q_{T\mid 0}\left(\cdot\mid\mathbf{x}_{0}\right)}\left[D_{\mathrm{KL}}\left(\mathbb{P}_{\mathbf{x}_{0}}\left(\cdot\mid\mathbf{x}_{t}\right)\|\mathbb{P}^{\theta}\left(\cdot\mid\mathbf{x}_{t}\right)\right)\right]-D_{\mathrm{KL}}\left(q_{T\mid 0}\left(\cdot\mid\mathbf{x}_{0}\right)\|\pi\right)\,$, (33)
which is equal to $\mathcal{L}_{1}$, according to (Lou et al., 2024).
3. Proving $K\geq-\mathcal{L}_{2}$
The derivation of $\mathcal{L}_{2}$ involves a second application of Jensen’s inequality on $K$:
$K$ $=\mathbb{E}_{q\left(\mathbf{x}_{0:T}\right)}\left[\sum_{t=1}^{T}\log\frac{p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}\right)}{q\left(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}\right)}+\log p\left(\mathbf{x}_{T}\right)\right]$ (34)
$=\mathbb{E}_{q\left(\mathbf{x}_{0:T}\right)}\left[\sum_{t=1}^{T}\log\left(\sum_{\tilde{\mathbf{x}}_{0}}q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\tilde{\mathbf{x}}_{0})\tilde{p}_{\theta}(\tilde{\mathbf{x}}_{0}\mid\mathbf{x}_{t})\right)\right]+C_{1}$
$\stackrel{{\scriptstyle(b)}}{{\geq}}\mathbb{E}_{q\left(\mathbf{x}_{0:T}\right)}\left[\sum_{t=1}^{T}\sum_{\tilde{\mathbf{x}}_{0}}q(\tilde{\mathbf{x}}_{0}\mid\mathbf{x}_{t-1})\log\left(\frac{q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})}{q(\tilde{\mathbf{x}}_{0}\mid\mathbf{x}_{t})}\tilde{p}_{\theta}(\tilde{\mathbf{x}}_{0}\mid\mathbf{x}_{t})\right)\right]+C_{1}$
$=\sum_{t=1}^{T}\mathbb{E}_{q(\mathbf{x}_{t},\mathbf{x}_{0})}[\log\tilde{p}_{\theta}(\mathbf{x}_{0}\mid\mathbf{x}_{t})]+C_{1}+C_{2}$
$=-\mathcal{L}_{2}\$
In summary, $\log p_{\theta}(\mathbf{x}_{0})\leq\mathcal{L}_{1}\leq\mathcal{L}_{2}$ holds.
Appendix F Implementation of the Resampler
We define a resampler $S:\mathbb{R}^{h}\rightarrow\mathbb{R}^{L\times D}$ , where $h$ is the length of the input vector, $L$ is the length of the sequence and $D$ is the hidden dimension of UniF2ace. Specifically, we define a learnable hidden latent matrix:
$\mathbf{M}_{0}\in\mathbb{R}^{L\times d},\quad\mathbf{M}_{0}=\text{LearnableParameter}$ (35)
where $d$ is the hidden dimension of the resampler. Its process involves:
1. Project the noise embedding $\mathbf{x}\in\mathbb{R}^{h}$ via
$\mathbf{H}=\mathbf{x}\mathbf{W}_{\text{in}}\in\mathbb{R}^{1\times d}$ (36)
2. Iteratively refine the latent matrix through $T$ layers, sucn as the $l$-th layer:
$\mathbf{M}^{\prime}_{l}$ $=\mathbf{M}_{l-1}+\text{MHA}\left(\mathbf{M}_{l-1},\,\text{Concat}(\mathbf{H},\mathbf{M}_{l-1})\right)$ (37)
$\mathbf{M}_{l}$ $=\mathbf{M}^{\prime}_{l}+\text{FFN}(\mathbf{M}^{\prime}_{l})$ (38)
where MHA denotes the Multi-Head Attention mechanism, FFN denotoes the Feed-Forward Network. In MHA, the query, key, and value are denoted as:
$Q_{l}$ $=M_{l-1}W_{Q}^{(l)}$ (39)
$K_{l}$ $=[H;M_{l-1}]W_{K}^{(l)}$ (40)
$V_{l}$ $=[H;M_{l-1}]W_{V}^{(l)}$ (41)
3. Project the final latent to the output space:
$\mathbf{Y}=\text{LayerNorm}(\mathbf{M}_{T}\mathbf{W}_{\text{out}})\in\mathbb{R}^{L\times D}$ (42)
This enables adaptive fusion of input vector into sequence features through learned latent queries.
Figure 8: More comparison of generated face images with other models. Fine-grained attributes are highlighted in the prompt.
Figure 9: More face images generated by UniF2ace
Figure 10: Comparison of captioning results and DeepSeeek-v3-based scores. We highlight fine-grained attributes with blue and errors in answers with red.
Figure 11: Prompts for building dataset. The first and second prompts are to GPT-4o, while the last prompt is to GPT-4. In the first prompt, the content in “[]” is used only when the image data includes built-in captions, such as in the MM-CelebA-HQ dataset.

### v5

@@
 1 Introduction
-Recently, unified multimodal models (UMMs) have emerged as a vibrant research area enabling cross-modality understanding and generation within a single “any-to-any” framework, marking a significant step toward artificial general intelligence (AGI) (Wu et al., 2024a; Shi et al., 2024; Li et al., 2024a; Zhou et al., 2024; Team, 2024; Xie et al., 2024).
+Recently, unified multimodal models (UMMs) have emerged as a vibrant research area enabling cross-modality understanding and generation within a single “any-to-any” framework, marking a significant step toward artificial general intelligence (AGI) (Wu et al., 2024a; Shi et al., 2024; Li et al., 2024a; Zhou et al., 2024a; Team, 2024; Xie et al., 2024; Zhou et al., 2026; 2024b).
 Given the central role of faces in daily life, applying this unified paradigm to achieve fine-grained face understanding and generation is essential for developing human-centric AGI.
@@
 3.2).
-Regarding the generation strategy, we recognize that the generation of fine-grained facial attributes is significantly more challenging than understanding tasks, as highlighted in prior studies (Du et al., 2017; Zhou et al., 2024; Xie et al., 2024).
+Regarding the generation strategy, we recognize that the generation of fine-grained facial attributes is significantly more challenging than understanding tasks, as highlighted in prior studies (Du et al., 2017; Zhou et al., 2024a; Xie et al., 2024).
 To address this, we harness the theory of score matching in discrete diffusion (Lou et al., 2024) and propose the dual discrete diffusion (D3Diff) training strategy, ensuring the meticulous synthesis of facial details.
-For network architecture, existing UMMs typically focus on dense architectures (Zhou et al., 2024; Xie et al., 2024) or solely on achieving token-level MoE (Deng et al., 2025), lacking the selective integration of instance features.
+For network architecture, existing UMMs typically focus on dense architectures (Zhou et al., 2024a; Xie et al., 2024) or solely on achieving token-level MoE (Deng et al., 2025), lacking the selective integration of instance features.
 To overcome these limitations, we introduce token-level and sequence-level MoE layers.
@@
 Dataset Face Resolution VQA Availability Image Caption Tokens (Avg.) Face Attributes (Avg.
-per Caption) LAION-Face (Zheng et al., 2022) Low 50M 16 2.7 FLIP-80M (Li et al., 2024c) 80M 22 4.4 FFHQ-Text (Zhou & Shimada, 2021) High 760 45 12.2 MM-CelebA-HQ (Karras et al., 2018) 30K 26 6.2 CelebV-Text (Yu et al., 2023b) 70K 80 4.3 UniF2aceD-1M (Ours) (1M) 130K 120 17.7
+per Caption) LAION-Face (Zheng et al., 2022) Low 50M 16 2.7 FLIP-80M (Li et al., 2024c) 80M 22 4.4 FFHQ-Text (Zhou and Shimada, 2021) High 760 45 12.2 MM-CelebA-HQ (Karras et al., 2018) 30K 26 6.2 CelebV-Text (Yu et al., 2023b) 70K 80 4.3 UniF2aceD-1M (Ours) (1M) 130K 120 17.7
 Figure 4: UniF2aceD-1M contains high-resolution facial images, the largest number of facial attributes, 130K fine-grained image-caption pairs and 1 million VQAs.
@@
 We rigorously evaluate UniF2ace’s performance across both generation and understanding tasks on our UniF2aceD-1M test set.
-To provide a comprehensive assessment and verify the generalizability of our method, we also conduct evaluations on other public benchmarks, including FFHQ-Text (Zhou & Shimada, 2021), MM-CelebA (Xia et al., 2021), and CelebV-Text (Yu et al., 2023b).
+To provide a comprehensive assessment and verify the generalizability of our method, we also conduct evaluations on other public benchmarks, including FFHQ-Text (Zhou and Shimada, 2021), MM-CelebA (Xia et al., 2021), and CelebV-Text (Yu et al., 2023b).
 For generation tasks, we used VQAscore to measure the relevance of generated images to captions, reporting results based on CLIP-FlanT5-11B (VQAscore-CF5) (Lin et al., 2024b) and LLaVA-v1.5-13B (VQAscore-LV) (Liu et al., 2024c) for robust assessment.
@@
 1), our 1.8B parameter UniF2ace sets a new state-of-the-art, outperforming all competing UMMs on key generation metrics including FID, VQA-score, and VLM-score.
-Furthermore, the model also demonstrates robust generalization, consistently achieving leading scores on public cross-facial datasets such as FFHQ-Text (Zhou & Shimada, 2021), MM-CelebA (Xia et al., 2021), and CelebV-Text (Yu et al., 2023a) (Tab.
+Furthermore, the model also demonstrates robust generalization, consistently achieving leading scores on public cross-facial datasets such as FFHQ-Text (Zhou and Shimada, 2021), MM-CelebA (Xia et al., 2021), and CelebV-Text (Yu et al., 2023a) (Tab.
 2).
@@
 Chameleon (Team, 2024) and Emu3 (Wang et al., 2024c) employ an ARM to simultaneously manage both tasks, highlighting the advantages of autoregressive models in multitask settings.
-Additionally, Transfusion (Zhou et al., 2024) and Show-o (Xie et al., 2024) combine a text ARM with a visual DM, enabling seamless integration of image understanding and generation.
+Additionally, Transfusion (Zhou et al., 2024a) and Show-o (Xie et al., 2024) combine a text ARM with a visual DM, enabling seamless integration of image understanding and generation.
 These studies have advanced the fusion of visual and text generation models, enhancing performance on multimodal tasks.